OpenStudy (anonymous):

by the method of mathematical induction prove 10^n+3(4^n+2)+5 is divisible by 9

OpenStudy (zzr0ck3r):

is that 4^(n+2)?

OpenStudy (zzr0ck3r):

if not then it is not even true for n = 1

OpenStudy (zzr0ck3r):

@meejoy ?

OpenStudy (zzr0ck3r):

let n = 1 10+3(6) + 5 = 33 9 does not divide 33

OpenStudy (anonymous):

Let n=1 \[LHS=10^1+3(4^{1+2})+5\] \[=10+192+5\] \[=207\] which is divisible by 9. Therefore true for n=1 Assume the formula is true for n=k \[\large 10^k+3(4^{k+2})+5=9m\] where \[\large m \epsilon \mathbb{R}\] For n=k+1 \[\large 10^{k+1}+3(4^{k+1+2})+5\] \[\large =10^{k+1}+3(4^{k+3})+5\]

OpenStudy (zzr0ck3r):

@meejoy ??

OpenStudy (anonymous):

no @zzr0ck3r You got the equation wrong...

OpenStudy (zzr0ck3r):

I see

OpenStudy (anonymous):

yes?? @zzr0ck3r ..

OpenStudy (anonymous):

@Azteck : are you done with your solution?

OpenStudy (anonymous):

\[\large =(10^k\times 10)+3(4^k \times 4^3)+5\] \[\large =(10^k \times 10)+3(64)(4^k)+5\]

OpenStudy (anonymous):

Not yet.

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

Forget the last line I previously did.

OpenStudy (anonymous):

\[\large =10(10^k+3(4^{k+2})+5)-45-18(4^{k+2})\]

OpenStudy (anonymous):

\[\large =10(10^k+3(4^{k+2})+5)-9[5+2(4^{k+2})]\] No matter what you put as k, for k+1 will be in the same form since you're subtracting a multiple of 9.

OpenStudy (anonymous):

Therefore you solved it by mathematical induction. YOu can write your concluding lines. I'm a bit tired now.

OpenStudy (anonymous):

=10(10k+3(4k+2)+5)−45−18(4k+2) where this came from? im a bit confuse

OpenStudy (anonymous):

I purposely rigged the equation so I could get the same form to your original equation.

OpenStudy (anonymous):

That is still equal to the previous line before that.

OpenStudy (anonymous):

I wanted to get rid of 10^k+1

OpenStudy (anonymous):

So I factorised it. But if you factorise 10 out just to get rid of 10^k+1, you would have to subtract some values for the middle and last terms to balance that out. Here's a basic example.

OpenStudy (anonymous):

\[5^{n+1}-3x+4=5(5^n -3x+4)+12x-16\]

OpenStudy (anonymous):

I wanted to get rid of the 1 in the n+1 so it can just be 1.

OpenStudy (anonymous):

If I took the 5 out, I would also affect the middle and last terms. So I have to do some rigging.

OpenStudy (anonymous):

so it can just be n**

OpenStudy (zzr0ck3r):

that was noe of you:)

OpenStudy (anonymous):

noe?

OpenStudy (zzr0ck3r):

nice*

OpenStudy (zzr0ck3r):

lol sorry

OpenStudy (anonymous):

No worries. And thanks mate.

OpenStudy (anonymous):

@Azteck : im sorry/... peace :)

OpenStudy (anonymous):

@azteck: im sorry...

OpenStudy (anonymous):

Okay, are you reading what I wrote for you?

OpenStudy (anonymous):

im reading it..

OpenStudy (anonymous):

okay im done.. and then??

OpenStudy (anonymous):

and then you can turn the [5+2(4^k+2)] into a pronumeral like b. and then state "where b is an integer"

OpenStudy (anonymous):

owww my g.. my head is aching >_<

OpenStudy (anonymous):

That's what mathematical induction does to you if you don't focus.

OpenStudy (anonymous):

Have you done that?

OpenStudy (anonymous):

i think u'll find this way easier

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

lol the substitution method. Our teacher was telling us to do it what I did. some other classes do it your way @Tushara

OpenStudy (anonymous):

meejoy.... there are many different ways to do induction questions... some ppl find some ways easier than other ways

OpenStudy (anonymous):

sorry about my typing. i'm typing with one hand. eating with my other hand.

OpenStudy (anonymous):

@Azteck i was taught both ways in highscool, its good to understand both ways

OpenStudy (anonymous):

@Tushara : thanks.. i appreciated it .. and also to you @Azteck ..

OpenStudy (anonymous):

Yeah. My friend was teaching it to me that way. I found it pretty easy to solve, but sometimes, I make more mistakes using that method somehow, so I use the other one. I might revert to your method.

OpenStudy (anonymous):

No worries.

OpenStudy (anonymous):

but i still dont understand.. huhu T_T

OpenStudy (anonymous):

Mathematical induction requires some "rigging" in order to get your required form.

OpenStudy (anonymous):

oh ok... there are so many different methods for induction... the more logic and explaining u do, the better u get at induction, cos u can start to think in many different ways... substituion is by far the easiest method

OpenStudy (anonymous):

You need to learn to RIG an expression so that you can get to where you want that expression to go to. That sentence sounds confusing, but it makes sense.

OpenStudy (anonymous):

And @Tushara pretty much hit the target. You need to be logical as well.

OpenStudy (anonymous):

that is why some questions say: do not use the substitution method

OpenStudy (sirm3d):

\[10^{k+1}+3(4^{k+1+2})+5\\=10(10^k)+3(4)(4^{k+2})+5\\=(9+1)10^k+3(3+1)(4^{k+2})+5\\=9(10^k)+(10^k)+3(3)(4^{k+2})+3(1)(4^{k+2})+5\\=9(10^k)+(10^k)+9(4^{k+2})+3(4^{k+2})+5\\=[9(10^k)+9(4^{k+2})]+[10^k+3(4^{k+2})+5]\\=9[10^k+4^{k+2}]+[9m]\\=9\left(10^k+4^{k+2}+m\right) \]